# Gaussian Mixture Models Based Augmentation Enhances GNN Generalization

### v2

@@ 3.1 Mathematical Formalism
Let $\ell$ be a classification loss function with $L_{Lip}$ as a Lipschitz constant such as $\ell(\cdot,\cdot)\in[0,1].$ Then, with a probability at least $1-\delta$ over the samples $\mathcal{D}_{train}$, we have,
$\mathbb{E}_{\mathcal{G}\sim G}\left[\ell(\mathcal{G},\hat{\theta}_{aug})\right]-\mathbb{E}_{\mathcal{G}\sim G}\left[\ell(\mathcal{G},\theta_{\star})\right]\leq 2\mathcal{R}(\ell_{aug})+5\sqrt{\frac{2\log(4/\delta)}{N}}+2L_{Lip}\mathbb{E}_{\mathcal{G}\sim G}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\left\|\mathcal{G}^{\lambda}-\mathcal{G}\right\|\right].$
@@
$\mathcal{R}(\ell_{aug})\leq\mathcal{R}(\ell)+\max_{n\in\{1,\ldots,N\}}L_{Lip}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\left\|\mathcal{G}_{n}^{\lambda}-\mathcal{G}_{n}\right\|\right].$
Theorem 3.1 relies on the assumption that the loss function is Lipschitz continuous. This assumption is realistic given that the input node features and graph structures in real-world datasets are typically bounded. Additionally, we can ensure that the loss function is bounded within $[0,1]$ by composing any standard classification loss with a strictly increasing function that maps values to the interval $[0,1]$. A direct implication of Theorem 3.1 is that if we chose the right data augmentation strategy $A_{\lambda}$ that minimizes the expected distance between original graphs and augmented ones $\mathbb{E}_{\mathcal{G}\sim G}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\left\|\mathcal{G}^{\lambda}-\mathcal{G}\right\|\right]$, we can guarantee with a high probability that the data augmentation decreases both the Rademacher complexity and the generalization risk. On the other hand, if the distance is large, we cannot guarantee that data augmentation will outperform the normal training setting.
The findings of Theorem 3.1 hold for all norms defined on the graph input space. Specifically, let us consider the graph space $(\mathbb{G},\lVert\cdot\rVert_{\mathbb{G}})$ and the feature space $(\mathbb{X},\lVert\cdot\rVert_{\mathbb{X}})$, where $\lVert\cdot\rVert_{\mathbb{G}}$ and $\lVert\cdot\rVert_{\mathbb{X}}$ denote the norms applied to the graph structure and features, respectively. Assuming a maximum number of nodes per graph, which is a realistic assumption for real-world data, the product space $\mathbb{G}\times\mathbb{X}$ is a finite-dimensional real vector space, and all the norms are equivalent. Thus, the choice of norm does not affect the theorem, as long as the Lipschitz constant is adjusted accordingly. Additional details and insights on the graph distance metrics can be found in Appendix G.
@@ 3.2 Proposed Approach
Based on the theoretical findings, it is crucial to employ a data augmentation technique that effectively controls the term $\mathbb{E}_{\mathcal{G}\sim G}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\left\|\mathcal{G}^{\lambda}-\mathcal{G}\right\|\right]$, to achieve stronger generalization guarantees. This consideration leads us to explore universal approximators, particularly GMMs, which are well-suited for this purpose, and can effectively approximate any data distribution, c.f. Theorem 3.2.
@@
(Goodfellow et al., 2016), Page 65. A Gaussian mixture model is a universal approximator of densities, in the sense that any smooth density can be approximated with any specific nonzero amount of error by a Gaussian mixture model with enough components.
@@
To achieve this, we first train a standard GNN on the graph classification task using the training set. Next, we obtain embeddings for all training graphs using the READOUT output, resulting in =$\mathcal{H}=\left\{h_{\mathcal{G}_{n}}~{}s.t.~{}\mathcal{G}_{n}\in\mathcal{D}_{train}\right\}$. These embeddings are used as the basis for generating augmented training graphs. We then partition the training set $\mathcal{D}_{train}$ by classes, such that $\mathcal{D}_{train}=\bigcup_{c}\mathcal{D}_{c}$ where $\mathcal{D}_{c}=\left\{\mathcal{G}_{n}\in\mathcal{D}_{train}~{},~{}y_{n}=c\right\}$. The objective is to learn new graph representations from these embeddings, and create augmented data for improved training.
We use the EM algorithm to learn the best-fitting GMM for the embeddings of each cluster $\mathcal{D}_{c}$, denoted as $\mathcal{H}_{c}=\left\{\mathbf{h}_{\mathcal{G}_{n}}~{}s.t.~{}\mathcal{G}_{n}\in\mathcal{D}_{c}\right\}$. The EM algorithm finds maximum likelihood estimates for each cluster $\mathcal{H}_{c}$. We first initialize the GMM distribution as in Eq. 2. Given a number of Gaussian distributions $K$, we specifically initialize the mean vector $\mu_{k}$, the covariance vector $\Sigma_{k}$, and the weight $\pi_{k}$ of each Gaussian distribution. The process then evolves iteratively: (i) Evaluate the posterior probabilities $\left\{\gamma_{ik}\right\}_{i,k}$, using the values of the mean vectors and covariance matrix (E-Step) Watanabe et al. (2010).
$\gamma_{ik}=\frac{\pi_{k}\mathcal{N}(x_{i}\mid\bm{\mu}_{k},\bm{\Sigma}_{k})}{\sum_{j=1}^{K}\pi_{j}\mathcal{N}(x_{i}\mid\bm{\mu}_{j},\bm{\Sigma}_{j})}.$
(ii) Estimate new parameters $\bm{\mu}_{k}$, $\bm{\Sigma}_{k}$ and $\pi_{k}$ with the updated values of $\left\{\gamma_{ik}\right\}_{i,k}$ (M-Step),
$\bm{\mu}_{k}$ $=\frac{\sum_{i=1}^{N}\gamma_{ik}x_{i}}{\sum_{i=1}^{N}\gamma_{ik}},~{}~{}~{}~{}\pi_{k}=\frac{1}{N}\sum_{i=1}^{N}\gamma_{ik},$
$\bm{\Sigma}_{k}$ $=\frac{\sum_{i=1}^{N}\gamma_{ik}(x_{i}-\bm{\mu}_{k})(x_{i}-\bm{\mu}_{k})^{\top}}{\sum_{i=1}^{N}\gamma_{ik}}.$
@@
Once a GMM distribution $p_{c}$ is fitted for each cluster $\mathcal{D}_{c}$, we use this GMM to generate new augmented data by sampling hidden representations from $p_{c}$. Each new sample drawn from $p_{c}$ is then assigned the corresponding cluster label $c$, ensuring that the augmented data inherits the label structure from the original clusters. After merging the hidden representations of both the original training data and the augmented graph data, we finetune the post-readout function, i.e., the final part of the GNN, which occurs after the readout function, on the graph classification task. Since the post-readout function consists of a linear layer followed by a Softmax function, the finetuning process is relatively fast. To evaluate our model during inference on test graphs, we input the test graphs into the GNN layers trained in the initial step to compute the hidden graph representations. For the post-readout function, we use the weights obtained from the second stage of training. Algorithm 1 and Figure 1 provide a summary of this approach.
@@
Figure 1: Illustration of GMM-GDA: Step 1. We first train the GNN on the graph classification task using the training graphs. Step 2. Next, we utilize the weights from the message passing layers to generate graph representations for the training graphs. Step 3. A GMM is then fit to these graph representations, from which we sample new graph representations. Step 4. Finally, we fine-tune the post-readout function for the graph classification task, using both the original training graphs and the augmented graph representations. For inference on the test set, we use the message passing weights trained in Step 1 and the post-readout function weights trained in Step 4.
Inputs: GNN of $T$ layers $f(\cdot,\theta)=\Psi\circ\texttt{READOUT}\left(\cup_{t=0}^{T}\{\texttt{AGGREGATE}^{(t)}\circ\texttt{COMBINE}^{(t)}(\cdot)\}\right)$
@@
            where $\Psi$ is the post-readout function, Graph classification dataset $\mathcal{D}$, Loss function $\mathcal{L},$
@@
  1. Train the GNN $f$ on the graph classification task on the training set $\mathcal{D}_{train}$;
  2. Use the trained Message Passing layers and the readout function to generate graph representation
             =$\mathcal{H}=\left\{\mathbf{h}_{\mathcal{G}_{n}}~{}s.t.~{}\mathcal{G}_{n}\in\mathcal{D}_{train}\right\}$ for the training set;
  3. Partition the training set $\mathcal{D}_{train}$ by classes, such that $\mathcal{D}_{train}=\bigcup_{c}\mathcal{D}_{c}$ where
            $\mathcal{D}_{c}=\left\{\mathcal{G}_{n}\in\mathcal{D}_{train}~{},~{}y_{n}=c\right\}$;
@@ 2. Use the trained Message Passing layers and the readout function to generate graph representation
      3.1. Fit a GMM distribution $p_{c}$ on the graph representations $\mathcal{H}_{c}=\left\{\mathbf{h}_{\mathcal{G}_{n}}~{}s.t.~{}\mathcal{G}_{n}\in\mathcal{D}_{c}\right\}$;
       3.2. Sample new graph representation $\widetilde{\mathcal{H}_{c}}=\{\widetilde{\mathbf{h}}~{}~{}s.t.~{}~{}\widetilde{\mathbf{h}}\sim p_{c}\}$ from the distribution $p_{c}$;
@@
  4. Finetune the post-Readout function $\Psi$ on the graph classification task directly on
            the new training set $\mathcal{H}=\cup_{c}\mathcal{H}_{c}$.
Algorithm 1 Detailed Steps in the GMM-GDA Algorithm
@@ 3.3 Time Complexity
One advantage of our approach is its efficiency, as it generates new augmented graph representations with minimal computational time. Unlike baseline methods, which apply augmentation strategies to each individual training graph (or pair of graphs in Mixup-based approaches) separately, our method learns the distribution of graph representations across the entire training dataset simultaneously using the EM algorithm (Ng, 2000). If $N=\left|\mathcal{D}_{train}\right|$ is the number of training graphs in the dataset, $d$ is the dimension of graph hidden representations $\left\{\mathbf{h}_{\mathcal{G}},~{}\mathcal{G}\in\mathcal{D}_{train}\right\}$, and $K$ is the number of Gaussian Components in the GMM, then the complexity to fit a GMM on $T$ iterations is $\mathcal{O}(N\cdot K\cdot T\cdot d^{2})$ (Yang et al., 2012). We compare the data augmentation times of our approach and the baselines in Table 7. Due to our different training scheme, i.e., where we first train the message passing layers and then train the pooling function after learning the GMM distribution, we measured the total backpropagation time and compared it with the backpropagation time of the baseline methods. The training time of baseline models varies depending on the augmentation strategy used, specifically, whether it involves pairs of graphs or individual graphs. Even in cases where a graph augmentation has a low computational cost for some baselines, training can still be time consuming as multiple augmented graphs are required to achieve satisfactory test accuracy. In contrast, GMM-GDA generates only one augmented graph per training graph, demonstrating effective generalization on the test set. Overall, our data augmentation approach is highly efficient during the sampling of augmented data, with minimal impact on the overall training time.

### v3

@@ 3.1 Formalism of Graph Data Augmentation
Let $\ell$ be a classification loss function with $\text{L}_{\text{Lip}}$ as a Lipschitz constant and $\ell(\cdot,\cdot)\in[0,1].$ Then, with a probability at least $1-\delta$ over the samples $\mathcal{D}_{\text{train}}$, we have,
$\begin{split}\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\ell(\mathcal{G},\hat{\theta}_{\text{aug}})\right]-\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\ell(\mathcal{G},\theta_{\star})\right]\leq 2\mathcal{R}(\ell_{\text{aug}})+\\ 5\sqrt{\frac{2\log(4/\delta)}{N}}+2\text{L}_{\text{Lip}}\mathbb{E}_{\mathcal{G}\sim\mathcal{D},\widetilde{\mathcal{G}}\sim A_{\lambda}}\left[\left\|\widetilde{\mathcal{G}}-\mathcal{G}\right\|\right].\end{split}$
@@
$\mathcal{R}(\ell_{\text{aug}})\leq\mathcal{R}(\ell)+\max_{n}\text{L}_{\text{Lip}}\mathbb{E}_{\widetilde{\mathcal{G}}_{n}^{m}\sim A_{\lambda}}\left[\left\|\widetilde{\mathcal{G}}_{n}^{m}-\mathcal{G}_{n}\right\|\right].$
Theorem 3.1 relies on the assumption that the loss function is Lipschitz continuous. This assumption is realistic, given that the input node features and graph structures in real-world datasets are typically bounded, i.e., node features are typically normalized or constrained within a fixed range, while graph structures, represented by adjacency matrices or their normalized forms, have bounded spectral properties, ensuring a constrained input space. Additionally, we can ensure that the loss function is bounded within $[0,1]$ by composing any standard classification loss with a strictly increasing function that maps values to the interval $[0,1]$. We provide the proof of this theorem in Appendix A.
The findings of Theorem 3.1 hold for all norms defined on the graph input space. Specifically, let us consider the graph structure space $(\mathbb{A},\lVert\cdot\rVert_{\mathbb{A}})$ and the feature space $(\mathbb{X},\lVert\cdot\rVert_{\mathbb{X}})$, where $\lVert\cdot\rVert_{\mathbb{A}}$ and $\lVert\cdot\rVert_{\mathbb{X}}$ denote the norms applied to the graph structure and features, respectively. Assuming a maximum number of nodes per graph, which is a realistic assumption for real-world data, the product space $\mathbb{A}\times\mathbb{X}$ is a finite-dimensional real vector space, and all the norms are equivalent. Thus, the choice of norm does not affect the theorem, as long as the Lipschitz constant is adjusted accordingly. Additional details on graph distance metrics and the comparison between the Lipschitz constants of GCN and GIN are provided in Appendices I and D.
The variety of distance metrics on the original graph space offers different upper bounds, which leads to distinct criteria for data augmentation based on these metrics to control the upper bound in Theorem 3.1. Instead, we propose to shift the focus to the hidden representation space of graphs, where we aim to derive more consistent and meaningful augmentations. If $\text{L}_{\text{Lip}}$ is taken to be the Lipschitz constant of the post-readout function only, then Theorem 3.1 still applies for the norm on the graph-level embeddings produced by the readout function, i.e., $\|\mathbf{h}_{\widetilde{\mathcal{G}}}-\mathbf{h}_{\mathcal{G}}\|$.
Working at the level of hidden representations of graphs, rather than directly in the graph input space, offers additional advantages. Hidden representations capture both the structural information and node features of each graph, enabling augmentation that enhances the generalization of both aspects simultaneously. Moreover, node alignment is needed to compare the original and augmented graphs, which is computationally expensive. By operating on hidden representations instead, node alignment becomes unnecessary. Furthermore, as we will discuss in Section 3.4, the effectiveness of augmented data depends on the specific GNN architecture. By leveraging graph representations learned through a GNN, we ensure that the augmentation process remains architecture-specific, aligning with the inductive biases of the chosen model.
@@ 3.2 Proposed Approach
Based on the theoretical findings, it is crucial to employ a data augmentation technique that effectively controls the term $\mathbb{E}_{\mathcal{G}\sim\mathcal{D},\widetilde{\mathcal{G}}\sim A_{\lambda}}\left[\|\mathbf{h}_{\widetilde{\mathcal{G}}}-\mathbf{h}_{\mathcal{G}}\|\right]$ which measures the expected deviation in the hidden representations of graphs under augmentation, to achieve stronger generalization guarantees. To better understand this term, we express it in terms of graph representations rather than graphs themselves,
$\mathbb{E}_{\mathcal{G}\sim\mathcal{D},\widetilde{\mathcal{G}}\sim A_{\lambda}}\left[\|\mathbf{h}_{\widetilde{\mathcal{G}}}-\mathbf{h}_{\mathcal{G}}\|\right]$ $=\mathbb{E}_{\mathbf{h}\sim\delta_{\mathcal{D}},\widetilde{\mathbf{h}}\sim Q_{\lambda}}\left[\|\widetilde{\mathbf{h}}-\mathbf{h}\|\right],$
where $Q_{\lambda}$ represents the augmentation strategy at the level of hidden representations, replacing $A_{\lambda}$, which operates at the graph level, and $\delta_{\mathcal{D}}:\mathbf{h}\mapsto\frac{1}{N}\sum_{n=1}^{N}\delta_{\mathbf{h}_{\mathcal{G}_{n}}}(\mathbf{h})$ represents the Dirac distribution over the training graph representations, capturing the empirical distribution of graph embeddings. In Proposition 3.2, we upper bound this expected perturbation.
Proposition 3.2.
Let $\delta_{\mathcal{D}}$ denote the discrete distribution of the training graph representations. Suppose we sample new augmented graph representations from a distribution $Q_{\lambda}$ defined on the support of $\delta_{\mathcal{D}}$. Then, the following inequality holds,
$\begin{split}\mathbb{E}_{\mathbf{h}\sim\delta_{\mathcal{D}},\widetilde{\mathbf{h}}\sim Q_{\lambda}}\left[\|\mathbf{h}-\widetilde{\mathbf{h}}\|\right]\leq~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\\ \sqrt{2}\cdot\sup_{\begin{subarray}{c}\mathbf{h}\sim\delta_{\mathcal{D}}\\ \widetilde{\mathbf{h}}\sim Q_{\lambda}\end{subarray}}\|\mathbf{h}-\widetilde{\mathbf{h}}\|\left(\sqrt{\mathcal{KL}(\delta_{\mathcal{D}}\parallel Q_{\lambda})}+\sqrt{2}\right),\end{split}$
where $\mathcal{KL}(\cdot\parallel\cdot)$ denotes the Kullback-Leibler divergence.
We provide a proof of Proposition 3.2 in Appendix B. A way to control the left side of this inequality is to choose a generator $Q_{\lambda}$ that minimizes both the $\mathcal{KL}(\delta_{\mathcal{D}}\parallel Q_{\lambda})$ and the supremum distance $\sup_{\mathbf{h}\sim\delta_{\mathcal{D}},\widetilde{\mathbf{h}}\sim Q_{\lambda}}\|\mathbf{h}-\widetilde{\mathbf{h}}\|$. Various universal approximators can be used to minimize $\mathcal{KL}(\delta_{\mathcal{D}}\parallel Q_{\lambda})$, including generative models like Generative Adversarial Networks (GANs) (Yang et al., 2012). These models are capable of approximating any probability distribution, making them powerful tools for learning complex augmentations. However, we specifically choose Gaussian Mixture Models (GMMs), which are well-suited for this purpose, and can effectively approximate any data distribution, c.f. Theorem 3.3. GMMs are computationally fast compared to other generative approaches, making them suitable for large-scale graph datasets. As shown in Appendix F, GMM-based augmentation yields better results compared to alternative generative strategies. Moreover, due to the exponential decay of Gaussian distributions, the supremum distance $\sup_{\mathbf{h}\sim\delta_{\mathcal{D}},\widetilde{\mathbf{h}}\sim Q_{\lambda}}\|\mathbf{h}-\widetilde{\mathbf{h}}\|$ is naturally constrained, ensuring a better control of the expected distance $\mathbb{E}_{\mathbf{h}\sim\delta_{\mathcal{D}},\widetilde{\mathbf{h}}\sim Q_{\lambda}}\left[\|\mathbf{h}-\widetilde{\mathbf{h}}\|\right]$.
@@
(Goodfellow et al., 2016, Page 65) A Gaussian mixture model is a universal approximator of densities, in the sense that any smooth density can be approximated with any specific nonzero amount of error by a Gaussian mixture model with enough components.
@@
To achieve this, we first train a standard GNN on the graph classification task using the training set. Next, we obtain embeddings for all training graphs using the READOUT output, resulting in =$\mathcal{H}=\left\{\mathbf{h}_{\mathcal{G}_{n}}~{}\text{s.t.}~{}\mathcal{G}_{n}\in\mathcal{D}_{\text{train}}\right\}$. These embeddings are used as the basis for generating augmented training graphs. We then partition the training set $\mathcal{D}_{\text{train}}$ by classes, such that $\mathcal{D}_{\text{train}}=\bigcup_{c}\mathcal{D}_{c}$ where $\mathcal{D}_{c}=\left\{\mathcal{G}_{n}\in\mathcal{D}_{\text{train}}~{},~{}y_{n}=c\right\}$. The objective is to learn new graph representations from these embeddings, and create augmented data for improved training.
We use the EM algorithm to learn the best-fitting GMM for the embeddings of each partition $\mathcal{D}_{c}$, denoted as $\mathcal{H}_{c}=\left\{\mathbf{h}_{\mathcal{G}_{n}}~{}\text{s.t.}~{}\mathcal{G}_{n}\in\mathcal{D}_{c}\right\}$. The EM algorithm finds maximum likelihood estimates for each cluster $\mathcal{H}_{c}$, following the procedure described in (Bishop & Nasrabadi, 2006).
@@
Once a GMM distribution $p_{c}$ is fitted for each partition $\mathcal{D}_{c}$, we use this GMM to generate new augmented data by sampling hidden representations from $p_{c}$. Each new sample drawn from $p_{c}$ is then assigned the corresponding partition label $c$, ensuring that the augmented data inherits the label structure from the original partitions. After merging the hidden representations of both the original training data and the augmented graph data, we fine-tune the post-readout function, i.e., the final part of the GNN, which occurs after the readout function, on the graph classification task. Since the post-readout function consists of a linear layer followed by a Softmax function, the finetuning process is relatively fast. To evaluate our model during inference on test graphs, we input the test graphs into the GNN layers trained in the initial step to compute the hidden graph representations. For the post-readout function, we use the weights obtained from the second stage of training. Algorithm 1 and Figure 1 provide a summary of the GRATIN model.
@@
Figure 1: Illustration of GRATIN. Step 1. We first train the GNN on the graph classification task using the training graphs. Step 2. Next, we utilize the weights from the message passing layers to generate graph representations for the training graphs. Step 3. A GMM is then fit to these graph representations, from which we sample new graph representations. Step 4. Finally, we fine-tune the post-readout function for the graph classification task, using both the original training graphs and the augmented graph representations. For inference on the test set, we use the message passing weights trained in Step 1 and the post-readout function weights trained in Step 4.
@@
Inputs: GNN of $T$ layers $f(\cdot,\theta)=\Psi\circ\texttt{READOUT}\circ g$, where $g$ is the composition of message passing layers, i.e, $g=\cup_{t=0}^{T}\{\texttt{AGGREGATE}^{(t)}\circ\texttt{COMBINE}^{(t)}(\cdot)\}$ and $\Psi$ is the post-readout function, graph classification dataset $\mathcal{D}$, loss function $\mathcal{L}$;
@@
1. Train GNN $f$ on the training set $\mathcal{D}_{\text{train}}$;
2. Use the trained message passing layers and the readout function to generate graph representations =$\mathcal{H}=\left\{\mathbf{h}_{\mathcal{G}_{n}}~{}s.t.~{}\mathcal{G}_{n}\in\mathcal{D}_{\text{train}}\right\}$ for the training set;
3. Partition the training set $\mathcal{D}_{\text{train}}$ by classes, such that $\mathcal{D}_{\text{train}}=\bigcup_{c}\mathcal{D}_{c}$ where $\mathcal{D}_{c}=\left\{\mathcal{G}_{n}\in\mathcal{D}_{\text{train}}~{},~{}y_{n}=c\right\}$;
@@ 1. Train GNN $f$ on the training set $\mathcal{D}_{\text{train}}$;
       3.1. Fit a GMM distribution $p_{c}$ on the graph representations $\mathcal{H}_{c}=\left\{\mathbf{h}_{\mathcal{G}_{n}}~{}\text{s.t.}~{}\mathcal{G}_{n}\in\mathcal{D}_{c}\right\}$;
3.2. Sample new graph representation $\widetilde{\mathcal{H}_{c}}=\{\widetilde{\mathbf{h}}~{}~{}\text{s.t.}~{}~{}\widetilde{\mathbf{h}}\sim p_{c}\}$ from the distribution $p_{c}$;
@@
4. Finetune the post-readout function $\Psi$ on the graph classification task directly on the new training set $\mathcal{H}=\cup_{c}\mathcal{H}_{c}$;
Algorithm 1 Graph classification with GRATIN
@@ 3.3 Time Complexity
One advantage of GRATIN is its efficiency, as it generates new augmented graph representations with minimal computational time. Unlike baseline methods, which apply augmentation strategies to each individual training graph (or pair of graphs in Mixup-based approaches) separately, our method learns the distribution of graph representations across the entire training dataset simultaneously using the EM algorithm (Ng, 2000). If $N=\left|\mathcal{D}_{\text{train}}\right|$ is the number of training graphs in the dataset, $d$ is the dimension of graph hidden representations $\left\{\mathbf{h}_{\mathcal{G}},~{}\mathcal{G}\in\mathcal{D}_{\text{train}}\right\}$, and $K$ is the number of Gaussian components in the GMM, then the complexity to fit a GMM on $T$ iterations is $\mathcal{O}(N\cdot K\cdot T\cdot d^{2})$ (Yang et al., 2012). We compare the data augmentation times of our approach and the baselines in Table 7. Due to our different training scheme, i.e., where we first train the message passing layers and then train the post-readout function after learning the GMM distribution, we have measured the total backpropagation time and compared it with the backpropagation time of the baseline methods. The training time of baseline models varies depending on the augmentation strategy used, specifically whether it involves pairs of graphs or individual graphs. Even in cases where a graph augmentation has a low computational cost for some baselines, training can still be time-consuming as multiple augmented graphs are required to achieve satisfactory test accuracy. In contrast, GRATIN generates only one augmented graph per training graph, demonstrating effective generalization on the test set. Overall, our data augmentation approach is highly efficient during the sampling of augmented data, with minimal impact on the training time. A complete analysis of the time complexity of GRATIN  and the baselines can be found in Appendix G.